# Effective Backdoor Mitigation Depends on the Pre-training Objective

### v3

@@ 3 Experiments
Table 1: This table shows the original Top-1 zero-shot Imagenet validation set accuracies and accuracy achieved after finetuning of the backdoored models. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$, i.e., the CleanCLIP approach. We choose the models with the highest accuracy which have an ASR value less than 5% (successful cleaning). The original ASR values for all models are more than 99%. Takeaway: The models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ maintain their original accuracy, while the ones trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ experience a huge drop relative to the starting accuracy ($\sim$20for model trained on CC3M dataset and 45% for model trained on CC6M dataset) after finetuning.
Pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ Pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$
Dataset Clean Data Size Orig. Acc. Clean Acc. (ASR $\leq$ 5%) $\uparrow$ Orig. Acc. Clean Acc. (ASR $\leq$ 5%) $\uparrow$
CC3M 100K 16.00% 16.49% 17.04% 14.16%
CC6M 100K 23.76% 24.04% 23.86% 13.05%
@@
In this section, we expound on the pre-training details for the models, following by their cleaning procedure, and the metrics we use to measure the performance.
Metrics
The models are evaluated for their Top-1 zero-shot accuracy on the Imagenet-1K validation set (referred to as Imagenet hereafter). Each of the 1,000 classes of Imagenet are described using sentences like: ‘a photo of a …’, ‘a tattoo of a …’, etc. We generate 80 such text templates for each class, and then pass them to the text encoder to produce an average text embedding for the class. During zero-shot classification, the prediction for an image is the class whose thus computed text embedding has the highest cosine similarity with the image embedding.
We also evaluate the attack success rate (ASR). In an apparent similarity to accuracy, the ASR of an attack for a backdoored model is defined as the percentage of triggered images that the model classifies as the target label. For measuring ASR, we add the trigger patch at random locations in all Imagenet validation set images and measure the percentage of them that are classified as “banana”. We measure both these metrics at the end of each finetuning epoch as any model encountered during the finetuning process is a good candidate for a cleaned model.
Pre-training
[…]
@@
Following CleanCLIP, we finetune the pre-trained models on a 100K, guaranteed to be clean, image-text pairs for 20 epochs using a batch size of 128 and AdamW as the optimizer. We perform extensive hyperparameter search and use various numbers of different learning rates (as many as 8 in some experiments and 14 in others, all with cosine scheduling and 50 warmup steps) for this process. Please refer to Appendix C for the set of learning rates explored for the cleaning procedure. Ideally, we would want to obtain a model that maintains the accuracy of the original backdoored model, while getting rid of its poison, i.e., very low ASR. We use three loss functions for finetuning:
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}$: CleanCLIP showed that finetuning with this loss did not change both the original model’s accuracy and ASR, and hence is an ineffective cleaning loss function. We reproduce these results for all models we trained.
@@
    $\mathcal{L}^{ft}_{\text{SSL}}$: CleanCLIP also showed that finetuning with this loss decreased the original model’s ASR at the expense of its accuracy, and hence is also an ineffective cleaning loss function. We also reproduce these results.
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$: CleanCLIP showed that finetuning with a combination of these losses decreased the original model’s ASR while not hurting its accuracy, and hence is an effective cleaning function. Our experiments show that while this observation is true for the models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$, it does not generalize to the models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
@@
Figure 1: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR, measured at the end of each finetuning epoch for the models pre-trained on the CC6M dataset. The finetuning is done with each of the three aforementioned losses. The red star in the top right corner corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: CleanCLIP successfully cleans the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left), while it fails for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
Figure 1 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each finetuning epoch for the models pre-trained on the CC6M dataset. Due to space limitations, we defer the plots for the CC3M dataset in Section D.1. For both the datasets, we observe that:
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$ serves as an effective cleaning loss for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left plot). The cleaned models maintain the accuracy of the original model but they have low ASR, which we consider successful cleaning. However, it does not lead to an effective cleaning of the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right plot). This model loses 45% of the original accuracy to obtain a low ASR ($\leq$ 5%).
Figure 7 in Section D.1 shows the metrics for the models trained on the CC3M dataset. The conclusion is similar to the model trained on CC6M dataset, i.e., the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model is successfully cleaned, however $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model lose about 20% of the original model’s accuracy to get a low ASR.
@@
For both datasets, our findings indicate that CleanCLIP is not effective in removing poison from the models pre-trained with a stronger pre-training objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, without a significant drop in accuracy. Table 1 gives the highest accuracy of the models which were successfully cleaned by CleanCLIP (ASR $\leq$ 5%).
Stopping Criterion for the Finetuning Trajectories
In the previous section, the ability to find a model with high accuracy and low ASR is considered a success for the CleanCLIP approach. However in practice, one would not be aware of the specific backdoor attack in the models, and therefore would not know which model to choose during the finetuning process, and when to stop the finetuning process. To address this practicality concern, we show the finetuning trajectories for the two pre-training objectives in Figure 2.
Figure 1(a) shows the finetuning trajectories of three runs with different learning rates for a model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ on the CC6M dataset. We observe that in all the three runs, the trajectory converges to a region of high accuracy and low ASR (top left corner), and the trajectories are smooth. This indicates a practitioner can finetune this model for as long as their resources allow, and choose the model at the end of the finetuning process, and is likely to obtain a model with high accuracy and low ASR.
Figure 1(b) shows the finetuning trajectory of three runs with different learning rates for a model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset. We observe that in all three runs, the model with the highest accuracy and an ASR $\leq$ 5% is an intermediate model in the finetuning trajectory and not the model at the end of the finetuning process. With continued finetuning, the model can both lose accuracy and see a gain in ASR, both of which are undesirable. Therefore, even if a cleaned model with decent accuracy and low ASR might exist for these models, for a practitioner, it is difficult to discern when to stop finetuning. Please refer to Appendix I for cleaning trajectories on all explored hyperparameters.
@@
Figure 2: Finetuning trajectories of models with different pre-training objectives. Successive epochs are shown with increasing size of the markers and intensity of the connecting line. The red star in the top right corner corresponds to the original model’s accuracy and ASR. Takeaway: Models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ converge to a region of high accuracy and low ASR as we continue to finetune. On the other hand, models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ fail to converge to a region of high accuracy and low ASR and continued finetuning can lead to both decreased accuracy and higher ASR. This makes determining the stopping criterion for finetuning process tricky.
Figure 3: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the models pre-trained on the CC6M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We measure accuracy and ASR at the end of each epoch. The red star in the top right corner corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: Even having 5 poisons in the cleaning dataset (i.e. 0.005% of the dataset, which is 10$\times$ cleaner than the pre-training data) hurts the cleaning process for both pre-training objectives, and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained models are hurt worse.
Cleaning under Non-ideal Conditions
@@
CleanCLIP assumes that the finetuning data is completely free of poisoned datapoints. In practice, this assumption can be violated even when considerable care is taken to ensure it. To simulate this real-world situation, we finetune models using cleaning data that has a few poisoned datapoints, specifically 5 and 10 poisoned datapoints out of 100K finetuning samples. Note that, these datasets are still 10$\times$ and 5$\times$ cleaner than the pre-training dataset respectively (in terms of the poisoning percentage). Figure 3 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each finetuning epoch for the models pre-trained on the CC6M dataset. We only show the finetuning with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We observe that having just 5 poisoned datapoints in the cleaning dataset severely weakens CleanCLIP for both the pre-training objectives. However for models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ alone, we found cleaned models that maintain the original model’s accuracy and achieve around 30-50% ASR.
@@
On the other hand, for the models pre-trained with the stronger objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, having just 5 poisoned examples renders the cleaning procedure completely ineffective. The models pre-trained on the CC6M dataset lose about 80% of the original model’s accuracy to obtain a low ASR ($\leq$ 5%) and no model has an ASR lower than 90% for the CC3M pre-trained model (Figure 8 in Section D.2).
@@
Our experiments highlight the fact that a stronger pre-training objective also affects the strength of poison induction, making the cleaning process difficult. Also, for a practitioner, when pre-training with a stronger objective, the decision of when to stop finetuning becomes non-trivial, as we show that the model at the end of the cleaning procedure is usually not the one with the lowest ASR and the best accuracy. The situation is further exacerbated when we even slightly relax the assumption of 100% poison-free cleaning data, which can be too stringent in practice.
4 Analysis of Stronger Pre-training Objective
We now perform several analysis experiments to understand the reason behind the difference in the poison removal ability of CleanCLIP across the two pre-training objectives. We also experiment with popular perturbation methods to attempt to remove the strongly induced poison. We refer the readers to Appendix C for hyperparameter details.

### v4

@@ 4 Experiments
In this section, we expound on the pre-training details for the models, followed by their cleaning procedure and the metrics we use to measure the model’s performance.
Metrics
The models are evaluated for their Top-1 zero-shot accuracy on the Imagenet-1K validation set (referred to as Imagenet hereafter). Each of the 1,000 classes of Imagenet is described using sentences like: ‘a photo of a …’, ‘a tattoo of a …’, etc. We generate 80 such text templates for each class (see Appendix C) and then pass them to the text encoder to produce an average text embedding for the class. During zero-shot classification, the prediction for an image is the class whose thus computed text embedding has the highest cosine similarity with the image embedding.
We also evaluate the attack success rate (ASR) of a model. In an apparent similarity to accuracy, the ASR of a backdoored model is defined as the percentage of triggered images that the model classifies as the adversary-chosen target label. For measuring ASR, we add the trigger patch at random locations in all Imagenet validation set images and measure the percentage of them that are classified as the target class, i.e., “banana”. We measure both these metrics at the end of each cleaning epoch as any model encountered during the cleaning process is a good candidate for a cleaned model.
Poison Induction by Training from Scratch
Table 2 shows the Top-1 zero-shot Imagenet validation set accuracy for the models trained from scratch using $\mathcal{L}^{pre}_{\text{MMCL}}$ and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on CC3M and CC6M datasets. For the smaller CC3M dataset, both the models achieve an accuracy of around 16–17%, and for the larger CC6M dataset, the models reach an accuracy of around 24%. Even though the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ attained higher accuracy than the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ alone, in order to have better visualization of the difference in performance of CleanCLIP on the two pre-training objectives, we deliberately choose models with similar starting accuracies. All the models, irrespective of the pre-training objective and the training dataset, reached more than 99% ASR (see Table 3 in Appendix A), implying that poisoning just 0.05% of the dataset is enough to attain very high ASR.
Table 2: This table shows the original Top-1 zero-shot Imagenet validation set accuracies and remaining accuracy after cleaning the models that were poisoned using BadNet by training from scratch. The cleaning is done using CleanCLIP, i.e., finetuning a poisoned model with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. For this table, we choose the models having the highest accuracy and ASR $\leq 5\%$ (successful cleaning). The original ASR values for all models are more than 99%. Takeaway: The models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ maintain their original accuracy after cleaning, while the ones trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ experience a huge drop relative to the starting accuracy ($\sim$20for model trained on CC3M dataset and 45% for model trained on CC6M dataset) after cleaning.
Trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ Trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$
Backdoor Dataset Clean Data Size Orig. Acc. Clean Acc. (ASR $\leq$ 5%) $\uparrow$ Orig. Acc. Clean Acc. (ASR $\leq$ 5%) $\uparrow$
BadNet CC3M 100K 16.00% 16.49% 17.04% 14.16%
BadNet CC6M 100K 23.76% 24.04% 23.86% 13.05%
@@
We clean the poisoned model by finetuning it on a 100K, guaranteed to be poison-free, image-text pairs for 20 epochs using a batch size of 128 and AdamW as the optimizer. We perform extensive hyperparameter search and use various learning rates (as many as 8 in some experiments and 14 in others, all with cosine scheduling and 50 warmup steps) for this process. Please refer to Appendix D for the set of learning rates explored for the cleaning procedure. Ideally, after cleaning we would want to obtain a model that maintains the accuracy of the original poisoned model, while getting rid of its poison, i.e., very low ASR. We use three different loss functions for the cleaning process:
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}$: CleanCLIP showed that finetuning with $\mathcal{L}^{ft}_{\text{MMCL}}$ did not change the original model’s accuracy and ASR, and hence is an ineffective cleaning loss. We reproduce these results for the models we trained.
@@
    $\mathcal{L}^{ft}_{\text{SSL}}$: CleanCLIP also showed that finetuning with $\mathcal{L}^{ft}_{\text{SSL}}$ decreased the original model’s ASR at the expense of its accuracy, and hence is also an ineffective cleaning loss. We also reproduce these results.
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$: CleanCLIP showed that finetuning with a combination of these losses decreased the original model’s ASR while not hurting its accuracy, and hence is an effective cleaning loss. Our experiments show that while this observation is true for the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$, however it does not generalize to the models trained with stronger pre-training objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$. This is our key finding.
@@
Figure 2 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each cleaning epoch for the models trained on the CC6M dataset. We defer the plots for the CC3M dataset in Section E.1 for space consideration. For both the datasets, we observe that:
@@
    $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$ serves as an effective cleaning loss for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left plot). The cleaned models maintain the accuracy of the original model, but they have low ASR, which we consider successful cleaning. However, it does not lead to an effective cleaning of the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right plot). Even the model that has the highest accuracy with a low ASR ($\leq$ 5%) is 45% less accurate than the original model, as shown in Figure 2.
Refer to caption 45% drop
Figure 2: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR, measured at the end of each cleaning epoch for the models trained on the CC6M dataset. The cleaning is done by finetuning the model with the three losses mentioned above. The red star in the top right corner (encircled in the black circle) corresponds to the model’s starting accuracy and ASR (before cleaning). For a successful cleaning, there should be models that maintain the model’s starting accuracy while having a low ASR (indicated by the red circle’s region in the top left). There are several models in the red circle in the left plot (successful clean), while there are no models in the red circle in the right plot (unsuccessful clean). Takeaway: CleanCLIP successfully cleans the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ (left), while it is ineffective for the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
@@
For both datasets, our findings indicate that CleanCLIP is not effective in removing poison from the models trained with a stronger pre-training objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, without a significant drop in accuracy. Table 2 gives the highest accuracy of the models which were successfully cleaned by CleanCLIP (ASR $\leq$ 5%).
Poison Induction by Finetuning a pre-trained model
We also induce poison by finetuning a pre-trained CLIP model (Radford et al., 2019). Concretely, we poison two models by finetuning them with $\mathcal{L}^{pre}_{\text{MMCL}}$ and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ respectively, using the CC6M dataset that had 3000 poisoned datapoints. We use a learning rate of $75e-$ 7 with AdamW optimizer with 10,000 warmup steps. After poisoning, these models achieve Top-1 zero-shot Imagenet set accuracy of $\sim 60\%$, much higher than the models trained from scratch (Figure 2). The ASR for the model poisoned with $\mathcal{L}^{pre}_{\text{MMCL}}$ is 99% and for the model poisoned with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ is 90%.
Findings from the Cleaning Procedure: We clean the poisoned models using CleanCLIP, i.e., further finetuning on a clean dataset with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. Figure 3 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each finetuning epoch for these two models. In this case, both the models experience a drop in accuracy to obtain a low ASR ($\leq$ 5%); however, the drop is much higher for the model when the poison was induced using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (33%, compared to a 17% drop for the model when the poison was induced using $\mathcal{L}^{pre}_{\text{MMCL}}$). This experiment corroborates our previous finding that CleanCLIP is less effective when the poison is induced using a combination of MMCL and SSL, irrespective of the fact whether the poison is induced via finetuning or by training from scratch.
Refer to caption 17% drop33% drop
Figure 3: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR, measured at the end of each cleaning epoch for the models poisoned by finetuning a CLIP pre-trained checkpoint on the CC6M dataset. The cleaning is done by finetuning the poisoned model with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. The red star in the top right corner (encircled in the black circle) corresponds to the original model’s accuracy and ASR (before cleaning). For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle in the top left). Takeaway: CleanCLIP is unable to successfully clean both the models; however, it performs much worse for the model poisoned with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
Poisoning a Different Backbone Architecture
For this experiment, we poison a model with a different backbone architecture, specifically ViTs. We poisoned these models by training them on the CC6M dataset with 3000 poisoned datapoints, and cleaned them using CleanCLIP. Figure 4 shows the cleaning plots for these models. We observe that similar to the previous experiments, the model trained with MMCL + SSL experiences a much larger drop in accuracy than the model just trained with MMCL, although in this case the model just trained with MMCL experiences a lot drop in accuracy as well.
Refer to caption 30% drop41% drop
Figure 4: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR, measured at the end of each cleaning epoch for the models trained on the CC6M dataset. The cleaning using CleanCLIP. The red star in the top right corner (encircled in the black circle) corresponds to the model’s starting accuracy and ASR (before cleaning). Takeaway: When a ViT backbone is poisoned, there are no cleaned models that maintain the original accuracies for both the pre-training losses, however the drop is much larger for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (right).
Poisoning Induction using a Different Poison
Due to space considerations, we present the results for the effectiveness of CleanCLIP when models are poisoned using label consistent backdoors in Appendix F. We observe that similar to the case of poisoning with BadNet, CleanCLIP is much less effective when the poison is induced using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ (16% loss in accuracy), compared to the case when it is induced using $\mathcal{L}^{pre}_{\text{MMCL}}$ (10% gain in accuracy).
@@
Figure 5: Finetuning trajectories of models with different pre-training objectives. Successive finetuning epochs are shown with increasing size of the markers and intensity of the connecting line. The red star in the top right corner (encircled in the black circle) corresponds to the original model’s accuracy and ASR. Takeaway: Models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ converge to a region of high accuracy and low ASR as we continue to finetune. On the other hand, models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ fail to converge to a region of high accuracy and low ASR, and continued finetuning can lead to both decreased accuracy and higher ASR. This makes determining the stopping criterion for the cleaning process for the latter models challenging.
Dependence of the Stopping Criterion on the Pre-training Objective
In the previous section, the ability to find a model with high accuracy and low ASR is considered a success for the CleanCLIP approach. However, in practice, one would not be aware of the ASR of the model being cleaned and, therefore, would not know when to stop the cleaning process. To highlight this practicality concern, we show the multiple cleaning trajectories for models trained using different pre-training objectives in Figure 5.
Figure 5(a) shows the trajectories of three cleaning runs with different learning rates for a model trained using $\mathcal{L}^{pre}_{\text{MMCL}}$ on the CC6M dataset. We observe that in all the three runs, the trajectory converges to a region of high accuracy and low ASR (top left corner), and the trajectories are smooth. This indicates that a practitioner can clean this model by finetuning the poisoned model for as long as their resources allow, and choose the model at the end of the finetuning process. They will likely obtain a model with high accuracy and low ASR, i.e., a successfully cleaned model.
Figure 5(b) shows the trajectories of three cleaning runs with different learning rates for a model trained using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset. We observe that in all three runs, the successfully cleaned model (high accuracy with a low ASR ($\leq$ 5%)) is an intermediate model in the trajectory, and not the model at the end of the process. With continued finetuning, the model can both lose accuracy and gain ASR, both of which are undesirable. Therefore, it is difficult for a practitioner to discern when to stop the finetuning process to obtain a clean model.
Therefore, the practicality of using CleanCLIP also depends on the pre-training objective of the model. Appendix K shows the cleaning trajectories for all explored hyperparameters.
Dependence of CleanCLIP on the Ideal Condition of the Dataset
@@
CleanCLIP assumes that the cleaning data is entirely free of poisoned datapoints. In practice, this assumption can be violated even when considerable care is taken to ensure it. To simulate this real-world situation, we clean models using data with a few poisoned datapoints, specifically 5 and 10 poisoned datapoints in the 100K cleaning datapoints. Note that these datasets are still, respectively, 10$\times$ and 5$\times$ cleaner than the original training dataset, illustrating a situation where the cleaning data is much cleaner than the training dataset but still not perfect.
Figure 6: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the cleaning process for the two models. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We measure accuracy and ASR at the end of each epoch. The red star in the top right corner (encircled in the black circle) corresponds to the original model’s accuracy and ASR. For a successful cleaning, there should be models that maintain the original model’s accuracy while having a low ASR (indicated by the red circle). Takeaway: Even having 5 poisons in the cleaning dataset (i.e. 0.005% of the dataset, which is 10$\times$ cleaner than the pre-training data) hurts the cleaning process for both pre-training objectives, and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ trained models are hurt worse.
Figure 6 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each cleaning epoch when two models trained from scratch on the CC6M dataset, one using $\mathcal{L}^{pre}_{\text{MMCL}}$ and the other using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ is cleaned by finetuning on this slightly poisoned dataset with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. We observe that having just 5 poisoned datapoints in the cleaning dataset severely weakens CleanCLIP for both the pre-training objectives.
@@
For models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$, we found cleaned models that maintain the original model’s accuracy and achieve around 30-50% ASR. On the other hand, for the models pre-trained with the stronger objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, having just 5 poisoned examples renders the cleaning procedure completely ineffective. The models pre-trained on the CC6M dataset lose about 80% of the original model’s accuracy to obtain a low ASR ($\leq$ 5%), and no model has an ASR lower than 90% for the CC3M pre-trained model (Figure 11).
@@
Our experiments highlight the fact that a stronger pre-training objective, like the combination of MMCL and SSL, also affects the strength of poison induction, making the cleaning process difficult. Also, for a practitioner, when pre-training with a stronger objective, the decision of when to stop finetuning becomes non-trivial, as we show that the model at the end of the cleaning procedure is usually not the one with the lowest ASR and the best accuracy. The situation is further exacerbated when we even slightly relax the assumption of 100% poison-free cleaning data, which can be too stringent in practice.
5 Analysis of the Stronger Pre-training Objective
We now perform several analysis experiments to understand the reason behind the difference in the poison removal ability of CleanCLIP across the two pre-training objectives. We also experiment with other methods to attempt to remove the poison when induced using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.